OpenStudy (anonymous):

A square has a side of 12 cm. if each side is increased in 15 %. what will be the new area of square?

OpenStudy (anonymous):

A= s^2 A= 12^2 A= 144 x .15?

OpenStudy (pradipgr817):

144*1.15*1.15

OpenStudy (jiteshmeghwal9):

as we know the side of a square are always equal

OpenStudy (anonymous):

1.15? or 0.15?

OpenStudy (jiteshmeghwal9):

what do u gt @Simonbane ???

OpenStudy (anonymous):

12.9?

OpenStudy (jiteshmeghwal9):

what do u gt ??

OpenStudy (anonymous):

51.6

OpenStudy (jiteshmeghwal9):

yes ! this is ur final answer :)

OpenStudy (anonymous):

can u explain again your drawing ?

OpenStudy (jiteshmeghwal9):

ohh !

OpenStudy (jiteshmeghwal9):

ok !

OpenStudy (jiteshmeghwal9):

\[\Huge{50\%\times 12}\]\[\Huge{\frac{3\cancel {15}}{5 \cancel {20} \cancel {100}} \times 3\cancel{12}}\]\[\Huge{9 \over 5}\] now add this by